OpenStudy (anonymous):

Solve: 8^x-1 = 2^x+1

OpenStudy (jkristia):

Do you have n idea how to start on this problem?

OpenStudy (lgbasallote):

hint: 8 = 2^3

OpenStudy (anonymous):

no i have no clue

OpenStudy (lgbasallote):

\[\large 8^{x-1} = 2^{x+1} \implies (2^3)^{x-1} = 2^{x+1}\]

OpenStudy (anonymous):

it wants me to solve for x though

OpenStudy (jkristia):

Since you have the same base (2) on both sides you can remove it, so you end up with 3(x+1) = x + 1, Now you can solve for x

OpenStudy (jkristia):

3(x-1), not 3(x+1) - it was a typo